# Kolmogorov-Arnold Network

### v2

@@ Abstract
In this work we propose $\mathbb{C}$KAN, a complex-valued KAN, to join the intrinsic interpretability of KANs and the advantages of Complex-Valued Neural Networks. We show how to transfer a KAN and the necessary associated mechanisms into the complex domain. To confirm that $\mathbb{C}$KAN meets expectations we conduct experiments on symbolic complex-valued function fitting and physically meaningful formulae as well as on a more realistic dataset from knot theory. Our proposed $\mathbb{C}$KAN is more stable and performs on par or better than real-valued KANs while requiring less parameters and a shallower network architecture, making it more explainable.
@@ I Introduction
In this work we aim to combine the advantages of KANs and CVNNs within our proposed Complex-Valued Kolmogorov-Arnold Network ($\mathbb{C}$KAN). By replacing the learnable real-valued functions on the edges of a KANs with learnable complex-valued functions we maintain the intrinsic explainability of KANs while also leveraging the direct use of complex-valued functions, making it more fit for problems requiring complex-valued calculations.
@@ II Related Work
II-A KANs
After the recent introduction of KANs by Liu et al. [1, 2] there has been a great variety of attempts to apply the ideas of KANs to different fields like image processing [11], satellite image segmentation [12], graph neural networks [13, 14] and even transformers [15]. Hou et al. [16] give a great overview of the different applications and extensions of KANs. Their explainability can be of great value in fields where machine learning approaches are strongly regulated like survival analysis in medicine or engineering [17]. Multiple works have already benchmarked the performance of KANs against MLPs [3, 18] and found KANs to be a more suitable alternative in some fields. Alter et al. [19] found that large-scale KANs are more robust against adversarial attacks as MLPs and thus form an interesting direction for further research in multiple fields.
@@ II-B Complex-Valued Neural Networks
After early introduction of CVNNs [22] they have lately risen in popularity since the introduction of building blocks for deep learning architectures [23]. Since then a lot of theoretical contributions have been made [24, 25, 26] to enable a multitude of applications [27, 28, 29, 30, 31].
The most closely related prior work in the complex domain are deep complex Radial Basis Function Neural Networks [32, 33]. However, in CRBFNs, the RBFs are applied to all inputs of a neuron (e.g. vertex of the computational graph) simultaneously, while in $\mathbb{C}$KANs the RBFs are applied on the edges of the computational graph to each value individually. Thus CRBFNs are architecturally more similar to classical MLPs with RBFs as activations functions, where we aim to adopt the KAN framework to the complex domain.
III Real-valued KANs
@@
III-A Real-valued RBFs
@@
To represent more complicated functions multiple RBFs are centered around uniformly distributed grid points $g_{i}$ inside a fixed grid with $i=0\dots G-1$ and $G$ denoting the number of grid points. All those RBFs are then combined in a weighted sum with learnable weights $w_{i}\in\mathbb{R}$ to construct one continuous function over the whole grid (3).
@@
IV Complex-valued KANs
To extend KANs into the complex-valued domain we need a way to learn functions $f:\mathbb{C}\to\mathbb{C}$ on every edge. Inspired by the use of RBFs in FastKAN we propose to construct those complex functions using multiple RBFs on a 2D grid. Furthermore we employ a complex-valued equivalent to the Sigmoid Linear Unit (SiLU) function used in KANs and make use of Batch Normalization to stay inside our fixed range grid. Finally we present a tool for visualizing the $\mathbb{C}$KAN to make it interpretable and explain its inner workings.
IV-A Complex-valued RBFs
@@
We can use this RBF in combination with complex-valued grid points and complex-valued weights to learn a complex-valued function. Let $w_{u,v}\in\mathbb{C}$ be learnable weights and $g_{u,v}\in\mathbb{C}$ complex-valued grid points on a grid with $G\times G$ grid points to cover the real and imaginary parts of the input.
@@ IV-B Residual Activation Function
In the real-valued KANs, a residual activation function is used to help the training of the univariate functions. To this end an activation function $\sigma$ is added to the sum of RBFs as explained above. Thus we chose to extend (5) to
@@
We propose to use $\mathbb{C}$SiLU as this residual activation function, a complex-valued equivalent to the SiLU-function (7) used in [1, 2, 7]:
@@
In our experiments in section V we study the difference in performance of both approaches.
@@ IV-C Complex-valued Batch Normalization
We employ these three approaches after every but the last layer in our $\mathbb{C}$KAN and study their performance in section V.
IV-D Interpretability of CVKAN
One major advantage of KANs is their inherent interpretability. Since the KAN learns on the edges, the learned univariate functions are directly applied to each input feature of the current layer. As proposed in [1, 2], we can utilize this property in two ways to explain the inner working of the KAN: We can give importance scores to all edges and vertices to understand which parts of the network significantly influence the output and we can visualize the learned functions on the edges to understand their input-output relation. Both of these can also be utilized for $\mathbb{C}$KANs.
@@ IV-D Interpretability of CVKAN
Fig. 4 presents an example of utilizing the standard deviation to assign relevance scores to edges and vertices as well as visualizing the individual learned edge functions.
For $\mathbb{C}$KAN, the inputs and outputs of edges are complex-valued. To visualize their behavior we thus have to visualize a function $\mathbb{C}\to\mathbb{C}$. We propose to visualize this as a colored 3D plot, where bases are the real and imaginary part of the input, the height is the magnitude of the output and the color shows the phase of the output. Due to the $2\pi$ periodicity of the phase, we choose a periodic color map. In Fig. 3 a small $\mathbb{C}$KAN is visualized as an example.
@@ V Experiments
To compare our complex-valued $\mathbb{C}$KAN with the real-valued KAN and FastKAN on complex-valued datasets we split the complex numbers of the input and output layer of real-valued KANs into real and imaginary parts. Thus we end up with real-valued KANs that have twice the input and output dimension of our $\mathbb{C}$KAN. We conduct three different experiments.
  • Just like in [1], we show that our network is capable of learning simple symbolic correlations between input and output based on synthetic formulae and outperforms its real-valued opponents on complex-valued problems.
@@
  • We use the knot dataset [37], which contains two complex-valued features and was also used by Liu et al. to study the original KAN [1] to analyze the performance of our $\mathbb{C}$KAN on a more realistic dataset and for classification.
For our experiments based on synthetic formulae we sample data points for each variable within our grid range and calculate the target output based on the symbolic formula. Our grid range is always $[-2,2]$ in $\mathbb{R}$ or $[(-2,-2i),(2,2i)]$ in $\mathbb{C}$ respectively. Because of Batch Normalization the resulting distribution has a standard deviation of $\sigma=1$ and mean $\mu=0$. We use a grid centered around zero spanning two standard deviations in each direction to optimize the trade off between a smaller grid size and minimizing the chance of the output of the former layer exceeding the grid. For the tabular knot dataset [37] we normalize each feature to be inside of our grid. If the dataset contains real-valued features we also feed them into our $\mathbb{C}$KAN as complex-valued numbers but set the imaginary part to zero.
@@ V-A Function Fitting
We chose four simple functions with one or two complex-valued input variables (cf. Table I) to demonstrate that our $\mathbb{C}$KAN is capable of learning basic correlations.
For each function from Table I one can determine the required model size by looking at the hierarchy of the function evaluation. Since we can learn any univariate function on a single edge, for $f_{1}$ a $1{\times}1$ $\mathbb{C}$KAN is optimal in theory. In the real-valued domain we need at least a $2{\times}3{\times}2$ KAN to represent the equivalent computation using only real numbers (cf. Table I). The function $f_{1\mathbb{R}}$ is an equivalent computation to $f_{1}$ using only real-valued numbers and twice as many inputs and outputs to accommodate the real and imaginary parts of the complex numbers separately. We need three neurons in the hidden layer to represent $x_{1}^{2}$, $x_{2}^{2}$ and $(x_{1}+x_{2})$. In the output layer we can then square the resulting value of the third hidden neuron and combine it with the (negative) identity of the first and second hidden neuron to construct the real and imaginary part of the output, each as a single real-valued number.
@@
As for the network size we chose the theoretically optimal architecture size for $\mathbb{C}$KAN and regular KAN respectively and also experimented with half the size of the optimal KAN architecture for $\mathbb{C}$KAN and vice versa double the optimal size of $\mathbb{C}$KAN for regular KAN. The grid size is $G=8$ per dimension for $\mathbb{C}$KAN and $G=8\cdot 8=64$ for real-valued KAN respectively.
@@
[…]
Figure 3: Visualization of a $\mathbb{C}$KAN for symbolic function fitting. Height of the graphs show magnitude of the learned function, colormap encodes the phase. Both functions on the lower edges represent quadratic functions with a constant offset. They sum to $z_{1}^{2}+z_{2}^{2}$ since their offsets cancel out, which can be see on the roots, which are identical besides a 90° rotation. The upper function represents a typical shape of a complex square function near the middle with some error artifacts on the edges of the grid.
Figure 4: Visualization of the first layer of a $\mathbb{C}$KAN on the knot dataset with the opacity of edges and nodes proportional to their relevance score.
@@
TABLE II: Results of $\mathbb{C}$KAN, FastKAN and KAN on four datasets for function fitting.
@@ V-B Physical Equations
The results of training on the holography dataset (cf. Table III) show that models with more parameters perform better. While our $\mathbb{C}$KAN has the best MSE score, FastKAN performs better than our approach on MAE with less trainable parameters. This hints that the $\mathbb{C}$KAN is more stable against outliers, but trades off this stability for a less accurate average result. When focusing on smaller - and thus more explainable - models, our model with size $3{\times}10{\times}1$ is the only small model to perform competitive with the best results.
A similar picture can be observed for the circuit dataset (cf. Table IV). In the MSE score, our $\mathbb{C}$KAN outperforms other methods, however in the MAE score FastKAN performs better. Notably, on this dataset scaling up the model does not increase the performance meaningfully, with the best overall MAE score being obtained with the smallest FastKAN model.
@@
TABLE III: Results on the Holography Dataset (15).
@@ V-C Knot Classification
The knot dataset [37] originates from knot theory. Davies et al. calculated 15 invariants on $\approx 240.000$ different knots, from which 13 are real-and two complex-valued. Thus we need a real-valued KAN with 17 input features or a $\mathbb{C}$KAN with 15 complex-valued input features, out of which 13 have their imaginary part set to zero. The task on this dataset is to classify the knots based on their invariants into 14 different classes representing their signature. Therefore the output layer’s width of both real- and complex-valued KANs is set to $14$ to produce a probability vector. Since these probabilities are real-valued we need to learn a function $\mathbb{C}\to\mathbb{R}$ in the output layer as described in subsection IV-A. The features of the dataset are normalized to fit in our grid of fixed size $[-2,2]$ for real-valued KANs and $[(-2-2i),(2+2i)]$ for $\mathbb{C}$KAN.
In Table V we list the results of KAN, FastKAN, $\mathbb{C}$KAN for hidden layer widths $\in\{1,2\}$ and a grid size of $G=8$. For the two real-valued KANs we also experimented with a grid size of $G=64$ each to conduct a fairer comparison of our $\mathbb{C}$KAN to the others. Our approach is the best-performing model on the knot dataset. While the best $\mathbb{C}$KAN also has the highest number of parameters, the $15{\times}1{\times}14$ $\mathbb{C}$KAN performs still better than FastKAN and approximately equally well as KAN with a similar amount of parameters. In Fig. 5 an exemplary confusion matrix of our best performing $\mathbb{C}$KAN is shown. The classes -12 and 12 are highly underrepresented in the dataset and thus $\mathbb{C}$KAN does not learn anything for these classes. Besides these outliers, all classes are learned well.
Fig. 4 shows the first layer of a $\mathbb{C}$KAN trained on the knot dataset with the opacity of nodes and edges showcasing the calculated relevance scores. Like in [1] the meridinal, here as one single complex number, and longitudinal translation are the most relevant features. To also quantitatively evaluate the explainability, we conduct a small study, where we exclude less relevant features and retrain the model on only the more relevant features. When reducing the dataset with this method to 7 or 3 features, we still obtain good results, only reducing the accuracy by $0.012$ and $0.028$ respectively (cf. Table VI). In contrast, when we leave out these 7 or 3 most relevant features, the performance drops to $\approx 30\%$. This shows that this method determines the relevance of features accurately.
@@
$\mathbb{C}$KAN $15{\times}1{\times}14$ 2921 $0.923\pm 0.001$ $0.212\pm 0.001$
@@
TABLE V: The best-performing $\mathbb{C}$KANs for hidden layer widths $\in\{1,2\}$, all experiments with FastKAN and KAN on the knot dataset.
Figure 5: Confusion matrix for $\mathbb{C}$KAN on the knot dataset normalized to probabilities within each row.

### v3

@@ Abstract
In this work we propose CVKAN, a complex-valued Kolmogorov-Arnold Network (KAN), to join the intrinsic interpretability of KANs and the advantages of Complex-Valued Neural Networks. We show how to transfer a KAN and the necessary associated mechanisms into the complex domain. To confirm that CVKAN meets expectations we conduct experiments on symbolic complex-valued function fitting and physically meaningful formulae as well as on a more realistic dataset from knot theory. Our proposed CVKAN is more stable and performs on par or better than real-valued KANs while requiring less parameters and a shallower network architecture, making it more explainable.
@@ I Introduction
In this work we aim to combine the advantages of KANs and CVNNs within our proposed Complex-Valued Kolmogorov-Arnold Network (CVKAN). By replacing the learnable real-valued functions on the edges of a KAN with learnable complex-valued functions we maintain the intrinsic explainability of KANs while also leveraging the direct use of complex-valued functions, making it more suitable for problems requiring complex-valued calculations.
@@ II Related Work
II-A Kolmogorov-Arnold Networks
After the recent introduction of KANs by Liu et al. [1, 2] there has been a great variety of attempts to apply the ideas of KANs to different fields like image processing [11], satellite image segmentation [12], graph neural networks [13, 14] and even transformers [15]. Hou et al. [16] give a detailed overview of the different applications and extensions of KANs. Their explainability can be of great value in fields where machine learning approaches are strongly regulated like survival analysis in medicine or engineering [17]. Multiple works have already benchmarked the performance of KANs against MLPs [3, 18] and found KANs to be a more suitable alternative in some fields. Alter et al. [19] found that large-scale KANs are more robust against adversarial attacks as MLPs and thus form an interesting direction for further research in multiple fields.
@@ II-B Complex-Valued Neural Networks
After early introduction of CVNNs [22, 23, 24, 25] they have lately risen in popularity since the introduction of building blocks for deep learning architectures [26]. Since then a lot of theoretical contributions have been made [27, 28, 29] to enable a multitude of applications [30, 31, 32, 33, 34].
The most closely related prior work in the complex domain are deep Complex Radial Basis Function Neural Networks [35, 36]. However, in CRBFNs, the RBFs are applied to all inputs of a neuron (e.g. vertex of the computational graph) simultaneously, while in CVKANs the RBFs are applied on the edges of the computational graph to each value individually. Thus CRBFNs are architecturally more similar to classical MLPs with RBFs as activation functions, where we instead aim to adopt the KAN framework to the complex domain.
III Real-Valued Kolmogorov-Arnold Networks
@@
III-A Real-Valued Radial Basis Functions
@@
To represent more complicated functions multiple RBFs are centered around uniformly distributed grid points $g_{i}$ inside a fixed grid with $i=0\dots G-1$ and $G$ denoting the number of grid points. All those RBFs are then combined in a weighted sum with learnable weights $w_{i}\in\mathbb{R}$ to construct one continuous function over the whole grid (3). In Fig. 1 we visualize such a function made up of $G=3$ RBFs.
@@
IV Complex-Valued Kolmogorov-Arnold Networks
To extend KANs into the complex-valued domain we need a way to learn functions $f:\mathbb{C}\to\mathbb{C}$ on every edge. Inspired by the use of RBFs in FastKAN we propose to construct those complex functions using multiple RBFs on a 2D grid. Furthermore we employ a complex-valued equivalent to the Sigmoid Linear Unit (SiLU) function used in KANs and make use of Batch Normalization to stay inside our fixed range grid. Finally we present a tool for visualizing the CVKAN to make it interpretable and explain its inner workings.
IV-A Complex-Valued Radial Basis Functions
@@
We can use this RBF in combination with complex-valued grid points and complex-valued weights to learn a complex-valued function. Let $w_{u,v}\in\mathbb{C}$ be learnable weights and $g_{u,v}\in\mathbb{C}$ complex-valued grid points on a grid with $G\times G$ grid points to cover the real and imaginary parts of the input. Then the complex-valued activation function $\Phi_{\mathbb{C}}:\mathbb{C}\to\mathbb{C}$ can be represented as:
@@ IV-B Residual Activation Function
In the real-valued KANs, a residual activation function is used to help the training of the univariate functions. To this end an activation function $\sigma$ is added to the sum of RBFs as explained above. Thus we chose to extend (5) to:
@@
We propose to use $\mathbb{C}$SiLU (8) as this residual activation function, a complex-valued equivalent to the SiLU-function (7) used in [1, 2, 7]:
@@
In our experiments in section V we analyze the difference in performance of both approaches.
@@ IV-C Complex-Valued Batch Normalization
We employ these three approaches after every but the last layer in our CVKAN and study their performance in section V.
IV-D Interpretability of Complex-Valued Kolmogorov-Arnold Networks
One major advantage of KANs is their inherent interpretability. Since the KAN learns on the edges, the learned univariate functions are directly applied to each input feature of the current layer. As proposed in [1, 2], we can utilize this property in two ways to explain the inner workings of the KAN: We can give importance scores to all edges and vertices to understand which parts of the network significantly influence the output and we can visualize the learned functions on the edges to understand their input-output relation. Both of these can also be utilized for CVKANs.
@@ IV-D Interpretability of Complex-Valued Kolmogorov-Arnold Networks
Fig. 4 presents an example of utilizing the standard deviation to assign relevance scores to edges and vertices as well as visualizing the most important learned edge function, i.e. the one with the highest relevance score.
For CVKAN, the inputs and outputs of edges are complex-valued. To visualize their behavior we thus have to visualize a function $\mathbb{C}\to\mathbb{C}$. We propose to visualize this as a colored 3D plot, where bases are the real and imaginary part of the input, the height is the magnitude of the output and the color shows the phase of the output. Due to the $2\pi$ periodicity of the phase, we choose a periodic color map. In Fig. 3 a small CVKAN is visualized as an example.
@@ V Experiments
To compare our complex-valued CVKAN with the real-valued KAN and FastKAN on complex-valued datasets we split the complex numbers of the input and output layer of real-valued KANs into real and imaginary parts. Thus we end up with real-valued KANs that have twice the input and output dimension of our CVKAN. We conduct three different experiments:
  • Just like in [1], we show that our network is capable of learning simple symbolic correlations between input and output based on synthetic formulae and outperforms its real-valued counterparts on complex-valued problems.
@@
  • We use the knot dataset [40], which contains two complex-valued features and was also used by Liu et al. to study the original KAN [1], to analyze the performance of our CVKAN on a more realistic dataset and for classification.
For our experiments based on synthetic formulae we sample data points for each variable within our grid range and calculate the target output based on the symbolic formula. Our grid range is always $[-2,2]$ in $\mathbb{R}$ or $[(-2,-2i),(2,2i)]$ in $\mathbb{C}$ respectively. Because of Batch Normalization the resulting distribution has a standard deviation of $\sigma=1$ and mean $\mu=0$. We use a grid centered around zero spanning two standard deviations in each direction to optimize the trade off between a smaller grid size and minimizing the chance of the output of the former layer exceeding the grid. For the tabular knot dataset [40] we normalize each feature to be inside of our grid. If the dataset contains real-valued features we also feed them into our CVKAN as complex-valued numbers but set the imaginary part to zero.
@@ V-A Function Fitting
We chose four simple functions with one or two complex-valued input variables (cf. Table I) to demonstrate that our CVKAN is capable of learning basic correlations.
For each function from Table I one can determine the required model size by looking at the hierarchy of the function evaluation. Since we can learn any univariate function on a single edge, for $f_{1}$ a $1{\times}1$ CVKAN is optimal in theory. In the real-valued domain we need at least a $2{\times}3{\times}2$ KAN to represent the equivalent computation using only real numbers (cf. Table I). The function $f_{1\mathbb{R}}$ is an equivalent computation to $f_{1}$ using only real-valued numbers and twice as many inputs and outputs to accommodate the real and imaginary parts of the complex numbers separately. We need three neurons in the hidden layer to represent $x_{1}^{2}$, $x_{2}^{2}$ and $(x_{1}+x_{2})$. In the output layer we can then square the resulting value of the third hidden neuron and combine it with the (negative) identity of the first and second hidden neuron to construct the real and imaginary part of the output, each as a single real-valued number.
@@
As for the network size we chose the theoretically optimal architecture size for CVKAN and regular KAN respectively and also experimented with half the size of the optimal KAN architecture for CVKAN and vice versa double the optimal size of CVKAN for regular KAN. The grid size is $G=8$ per dimension for CVKAN and $G=8\cdot 8=64$ for real-valued KAN respectively.
@@
The results of training on toy datasets generated by simple complex-valued formulae (cf. Table I) can be seen in Table II. For $z^{2}$ and $(z_{1}^{2}+z_{2}^{2})^{2}$ our CVKAN is by far the best performing model when compared to FastKAN and KAN with similar or even bigger numbers of parameters. For $z_{1}\cdot z_{2}$ CVKAN is slightly better than FastKAN while CVKAN requires only half the number of parameters. For $\sin(z)$ our approach is slightly worse than FastKAN, however FastKAN requires a bigger network size and thus $\approx 16$ times as many parameters to perform marginally better than CVKAN. With CVKAN we are also able to achieve on average lower standard deviations on the results of the 5-fold cross validation, emphasizing the stability of our proposed model.
Figure 3: Visualization of a CVKAN for symbolic function fitting. Height of the graphs show magnitude of the learned function, colormap encodes the phase. Both functions on the lower edges represent quadratic functions with a constant offset. They sum to $z_{1}^{2}+z_{2}^{2}$ since their offsets cancel out, which can be seen on the roots, which are identical besides a 90° rotation. The upper function represents a typical shape of a complex square function near the middle with some error artifacts on the edges of the grid.
Figure 4: Visualization of the first layer of a CVKAN on the knot dataset with the opacity of edges and nodes proportional to their relevance score. For better readability only the function corresponding to the most relevant blue input node for feature merid_translat_c is shown. In our visualization tool the individual functions and edges can be interactively selected by double-clicking on the corresponding nodes or by specifying the edge index $(l,i,j)$ to view only the function of edge $E_{l,j,i}$ in detail.
TABLE II: Results of CVKAN, FastKAN and KAN on four datasets for function fitting.
@@ V-B Physical Equations
The results of training on the holography dataset (cf. Table III) show that models with more parameters perform better. While our CVKAN has the best MSE score, FastKAN performs better than our approach on MAE with less trainable parameters. This hints that the CVKAN is more stable against outliers, but trades off this stability for a less accurate average result. When focusing on smaller - and thus more explainable - models, our model with size $3{\times}10{\times}1$ is the only small model to perform competitive with the best results.
A similar picture can be observed for the circuit dataset (cf. Table IV). In the MSE score, our CVKAN outperforms other methods, however in the MAE score FastKAN performs better. Notably, on this dataset scaling up the model does not increase the performance meaningfully, with the best overall MAE score being obtained with the smallest FastKAN model.
TABLE III: Results on the Holography Dataset (15).
@@
TABLE IV: Results on the Circuit Dataset (16).
@@ V-C Knot Classification
The knot dataset [40] originates from knot theory. Davies et al. calculated 15 invariants on $\approx 240.000$ different knots, from which 13 are real-and two complex-valued. Thus we need a real-valued KAN with 17 input features or a CVKAN with 15 complex-valued input features, out of which 13 have their imaginary part set to zero. The task on this dataset is to classify the knots based on their invariants into 14 different classes representing their signature. Therefore the output layer’s width of both real- and complex-valued KANs is set to $14$ to produce a probability vector. Since these probabilities are real-valued we need to learn a function $\mathbb{C}\to\mathbb{R}$ in the output layer as described in subsection IV-A. The features of the dataset are normalized to fit in our grid of fixed size $[-2,2]$ for real-valued KANs and $[(-2-2i),(2+2i)]$ for CVKAN.
In Table V we list the results of KAN, FastKAN and CVKAN for hidden layer widths $\in\{1,2\}$ and a grid size of $G=8$. For the two real-valued KANs we also experimented with a grid size of $G=64$ each to conduct a fairer comparison of our CVKAN to the others. Our approach is the best-performing model on the knot dataset. While the best CVKAN also has the highest number of parameters, the $15{\times}1{\times}14$ CVKAN performs still better than FastKAN and approximately equally well as KAN with a similar amount of parameters. In Fig. 5 an exemplary confusion matrix of our best performing CVKAN is shown. The classes -12 and 12 are highly underrepresented in the dataset and thus CVKAN does not learn anything for these classes. Besides these outliers, all classes are learned well.
Fig. 4 shows the first layer of a CVKAN trained on the knot dataset with the opacity of nodes and edges showcasing the calculated relevance scores. Like in [1] the meridinal, here as one single complex number, and longitudinal translation are the most relevant features. To also quantitatively evaluate the explainability, we conduct a small study, where we exclude less relevant features and retrain the model on only the more relevant features. When reducing the dataset with this method to 7 or 3 features, we still obtain good results, only reducing the accuracy by $0.012$ and $0.028$ respectively (cf. Table VI). In contrast, when we leave out these 7 or 3 most relevant features, the performance drops to $\approx 30\%$. This shows that this method determines the relevance of features accurately.
TABLE V: The best-performing CVKANs for hidden layer widths $\in\{1,2\}$, all experiments with FastKAN and KAN on the knot dataset.
@@
CVKAN $15{\times}1{\times}14$ 2921 $0.923\pm 0.001$ $0.212\pm 0.001$
@@
Figure 5: Confusion matrix for CVKAN on the knot dataset normalized to probabilities within each row.